A force of P N acting on a body of 5kg traveling with a velocity of 5ms-1 increases it velocity to 10ms-1. if the force acts in the direction of motion for a period of 5s, what is the value of P?


1
Expert's answer
2022-02-22T08:52:39-0500

Given:

"m=5\:\rm kg"

"v_1=5\:\rm m/s"

"v_2=10\:\rm m/s"

"\Delta t=5\:\rm s"


The Newton's second law says

"F=m\frac{\Delta v}{\Delta t}=5*\frac{10-5}{5}=5\:\rm N"